OpenStudy (jiteshmeghwal9):

By selling two chairs for Rs.1200 each, a man gains 10% on one chair & loses 10% on the other . Find his gain or loss per cent in the whole transaction.

OpenStudy (jiteshmeghwal9):

@satellite73 @Zarkon Plz help :)

OpenStudy (anonymous):

He gains 120, but as well as losses 120

OpenStudy (anonymous):

i guess we can do it with the numbers given. so the 1200 is a ten percent gain, meaning if i read it correctly that it is 110% what the chair cost the merchant the merchant therefore pain \(1200\div1.1=1090.91\) rounded for the chair, and made a profit of \(\$109.09\)

OpenStudy (jiteshmeghwal9):

ya i also did something like that but not getting the concept :(

OpenStudy (anonymous):

similarly the other 12000 represented at 10% loss, so he sold it for 90% of his cost his cost was therefore \(1200\div .9=1333.33\) rounded and he lost 133.33

OpenStudy (jiteshmeghwal9):

hmm... i solved this through right here only

OpenStudy (jiteshmeghwal9):

what is the next step ?

OpenStudy (anonymous):

then the net loss is \(109.09-133.33=-24.24\)

OpenStudy (jiteshmeghwal9):

yes !

OpenStudy (anonymous):

as a percent of the entire transaction it is \[-\frac{24.24}{2400}\]

OpenStudy (jiteshmeghwal9):

ok

OpenStudy (anonymous):

usually it is easier to work with numbers but this time it might be easier to work with variables

OpenStudy (jiteshmeghwal9):

yup!

OpenStudy (anonymous):

\[\frac{\frac{x}{1.1}-\frac{x}{.9}}{2x}\]

OpenStudy (anonymous):

oh wait that is not right, let me try again

OpenStudy (jiteshmeghwal9):

okay!

OpenStudy (anonymous):

if we replace \(1200\) by \(x\) it looks like we computed \[\frac{\frac{x}{1.1}-x+\frac{x}{.9}-x}{2x}\]

OpenStudy (anonymous):

now maybe some algebra to clean it up looks like we get \[\frac{1}{99}\]

OpenStudy (jiteshmeghwal9):

why ?

OpenStudy (anonymous):

actually it is the other way around \[\frac{x-\frac{x}{1.1}+x-\frac{x}{.9}}{2x}\]

OpenStudy (anonymous):

why as in how did i get \(\frac{1}{99}\)?

OpenStudy (anonymous):

multiply top and bottom by either 9.9 or 99 to clear the decimal, combine like terms etc and you will get it \(\frac{1}{99}=.010101...\) as a prercent it is a loss of \(1\%\) rounded, but really \(1\tfrac{1}{99}\%\)

OpenStudy (jiteshmeghwal9):

hmmmmm.. I see thanx :)

OpenStudy (anonymous):

yw